OpenStudy (anonymous):

The y-intercept of the line whose equation is 2x + 5y = 8 is i got 8?

OpenStudy (anonymous):

gotta solve for y. 5y=8-2x =y=(8/5)-2x/5 so y intercept should be 8/5

OpenStudy (anonymous):

Moldy85 has the right answer, For your reference, I posted the answer at http://www.oojih.com/show/algebra/linearfunction/ scroll down to the bottom to see the solution.

OpenStudy (anonymous):

y intercept is going to be the constant (no variable) term if you solve for y

OpenStudy (anonymous):

or y intercept means x =0